OpenStudy (anonymous):

Find side B of a right triangle if side A is 12 yd and side B is 15 yd.

OpenStudy (anonymous):

find side C...? do you mean if so pathygren therom....* A^2+B^2=C^2 plug it in and thats the answer

OpenStudy (anonymous):

So 12^2+ 15^2= C^2 add the two square roots then square root your answer and c^2 to get ___=c

OpenStudy (anonymous):

Thanks.